On the kickoff, the football was in the air for 3 seconds before hitting the ground 68 meters away. How fast did the ball leave the foot of the kicker?
1
Expert's answer
2020-01-27T09:23:08-0500

Time of flight of the football="3=\dfrac{2u\sin\theta}{g}-----(i)"


"68=u\cos\theta \times T"


"68=u \cos\theta \times 3"


"u=\dfrac{68}{3\cos\theta}--------(ii)"


From the equation (i) and (ii)


"3=\dfrac{2\times 68 \sin\theta}{3\cos\theta}"


"\tan \theta=\dfrac{9}{138}"


"\theta=3.731"



"68=u\cos (3.7) \times 3"

"u=\dfrac{68}{3\cos{3.7}} =\dfrac{68}{0.99\times 3}=22.89m/sec"